OpenStudy (tabdo1226):

What is the value of the first quartile (Q1) of the data set? Table with key of three one means thirty one

OpenStudy (tabdo1226):

@texaschic101

OpenStudy (tabdo1226):

@ILUVSOCCER

OpenStudy (tabdo1226):

will fan and medal

OpenStudy (iluvsoccer):

Sorry, i forgot this.

OpenStudy (tabdo1226):

@brynndelyn

OpenStudy (tabdo1226):

@ILUVSOCCER can you call some body that knows this

OpenStudy (iluvsoccer):

@Miracrown

OpenStudy (tabdo1226):

are you done @texaschic101

OpenStudy (tabdo1226):

is 12

OpenStudy (tabdo1226):

* is it 12

OpenStudy (tabdo1226):

i think it's 12

OpenStudy (texaschic101):

put your numbers down.. 45,51,52,58,61,63,66,72,74 Q2 (middle number) = 61 Q1 = (51 + 52)/2 = 103/2 = 51.5 Q3 = (66 + 72)/2 = 138/2 = 69

OpenStudy (texaschic101):

questions ?

OpenStudy (tabdo1226):

is it 69

OpenStudy (texaschic101):

no....Q1 = 51.5 Q1 can be found by first finding Q2 (the middle number)...then taking the average of all the numbers before Q2.

OpenStudy (tabdo1226):

so whats the answer

OpenStudy (texaschic101):

oops..not the average...the middle number of all the numbers before Q2 answer is Q1 = 51.5

OpenStudy (tabdo1226):

thank you

OpenStudy (texaschic101):

no problem :)

OpenStudy (tabdo1226):

can you help me with some more

OpenStudy (texaschic101):

I have a little more time....I can try